1. The force required to draw back the string of an archers bow by a distance 'x' from it equilibrium position is given by f=kx2 if all the energy stored in the bow is communicated to the arrow, show that the maximum range of the arrow is given by R=2kx2/3mg
1
Expert's answer
2021-09-29T17:49:26-0400

"F=kx^2,"

"A=\int Fdx=\frac{kx^3}3,"

"E=\frac{mv^2}2,"

"A=E,\implies v^2=\frac{2kx^3}{3m},"

"l_{max}=\frac{v^2}g=\frac{2kx^3}{3mg}=R."